Suppose that Maria has 130 coins consisting of pennies, nickels, and dimes. The number of nickels she has is 40 less than twice the number of pennies; the number of dimes she has is 10 less than three times the number of pennies. How many coins of each kind does she have?

Answer by lwsshak3(11628)

Suppose that Maria has 130 coins consisting of pennies, nickels, and dimes. The number of nickels she has is 40 less than twice the number of pennies; the number of dimes she has is 10 less than three times the number of pennies. How many coins of each kind does she have?
***
let x=number of pennies
2x-40=number of nickels
3x-10=number of dimes
...
x+(2x-40)+(3x-10)=130
x+2x-40+3x-10=130
6x=180
x=30
2x-40=60-40=20
3x-10=90-10=80
number of pennies=30
number of nickels=20
number of dimes=80